# HOTS3D: Hyper-Spherical Optimal Transport for Semantic Alignment of Text-to-3D Generation

### v1

@@ 4.2.3 Qualitative Results
The qualitative comparison is presented in Fig. 5, indicating that the baseline methods [2, 18, 8, 33] encounter challenges in generating precise and realistic 3D objects, resulting in semantic inconsistency or incomplete 3D shapes. The results of CLIP-Forge exhibit significant grid artifacts and have low consistency with the text. PureCLIPNeRF can produce 3D content with richer textures and colors, significantly improving text consistency, but has poor control over the geometric shape of the target object. TAPS3D tends to generate simple-styled tables and chairs, while Shap·E lacks control over the geometric shape of objects, both of them need to improve their semantic consistency. SDFusion’s results are more reasonable than the aforementioned baselines, but there is also an obvious issue of text and shape mismatch. In contrast, our method generated more semantically coherent and higher-quality 3D objects by incorporating the SOT map to align text features and shape features. For instance, for the “U shaped legs” in the first row, for the “2 tables hooked together” in the third row, and the “mix rexine material” in the fifth row, the 3D shapes generated by our model match the text very well. In addition, our results are better in terms of detail and overall quality.
4.2.4 Visual explanation of SOT
To demonstrate the effectiveness of the proposed SOT algorithm in section 3.2, we conducted validation experiments on synthesized spherical data, and the visualization results are shown in Fig. 6. Here, we employed Mollweide projection[78] and vertical projection to visualize the results. The Mollweide projection is an area-preserving [79] projection generally used for maps of the world or celestial sphere. It trades accuracy of angle and shape for accuracy of proportions in area, making it useful for maps depicting global distributions. Specifically, in the first three columns of Fig. 6, we used Mollweide projection, and in the last column, we employed vertical projection to provide different viewing perspectives. In the first row of Fig. 6, the source domain data $\boldsymbol{X}$ follows four modes Von Mises-Fisher distribution (vMF) [80], while the target data $\boldsymbol{Y}$ follows uniform distribution on two discontinuous one-eighth spheres. In the second row, the source $\boldsymbol{X}$ follows single mode Von Mises-Fisher distribution and the target $\boldsymbol{Y}$ follows six modes Von Mises-Fisher distribution.
The vMF distribution is a distribution on $S^{d-1}$ characterized by a concentration parameter $\kappa>0$ and a location parameter $\mu\in S^{d-1}$ through the density
$\forall\theta\in S^{d-1},\ f_{\mathrm{vMF}}(\theta;\mu,\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}\exp(\kappa\mu^{T}\theta),$ (18)
where $I_{\nu}(\kappa)=\frac{1}{2\pi}\int_{0}^{\pi}\exp(\kappa\cos(\theta))\cos(\nu\theta)\mathrm{d}\theta$ is the modified Bessel function of the first kind. It is often referred to as the analogous of the Gaussian distribution on the hyper-sphere[81].
Specifically, we set the mean direction ($\mu$ in Euclidean space) to $(1,0,0)$, $(0,1,0)$, $(-1,0,0)$ and $(0,-1,0)$ respectively, and then adopt commonly used open source code [^1] https://github.com/dlwhittenbury/von-Mises-Fisher-Sampling/ to generate a mixed vMF distribution with 4 modes. For the target data $Y$, we uniformly sampled on two discontinuous one-eighth spheres, whose elevation $\theta$ and azimuth $\varphi$ ranges are $\frac{\pi}{3}\leq\theta\leq\frac{5\pi}{6},0\leq\varphi\leq\frac{\pi}{2}$ and $\frac{\pi}{3}\leq\theta\leq\frac{5\pi}{6},\frac{\pi}{2}+\frac{\pi}{12}\leq\varphi\leq\pi+\frac{\pi}{12}$, respectively.
For another experiment shown in the second row in Fig. 6, we set the mean direction to $(1,0,0)$, $(0,1,0)$, $(0,0,1)$, $(-1,0,0)$, $(0,-1,0)$, and $(0,0,-1)$ respectively to generate a mixed vMF distribution with 6 modes, and take it as the target distribution. Then, we set the mean direction as $(1,0,0)$ to generate a single mode vMF distribution and take it as the source distribution.
From the results on the right, it can be seen that the distribution of $T(\boldsymbol{X})$ obtained from our SOT map can align well with the target distribution of $\boldsymbol{Y}$. This demonstrates that the proposed SOT map not only effectively maintains clear boundaries between discontinuous regions, but also approximates target distributions with intricate support.
4.2.5 Diversity of generated contents
The proposed HOTS3D has a notable level of creativity. As illustrated in Fig. 7, it can generate varied 3D models even with the same text prompt. In the first row of Fig. 7, the first two generated tables exhibit significant differences in leg design, yet both effectively maintain consistency with the text prompt. As the text on the top right does not specify strict constraints on the shape, our HOTS3D creatively explores the geometric shapes of both the tabletop and table legs. For the chairs in the second row, our model innovatively designs the armrests and backrests while ensuring alignment with the textual description. For the fourth text, the results generated by our model not only meet the descriptions like “wooden material”, “thick rectangular top” and “four small legs”, but also present diverse combinations of tabletops and legs. Moreover, as shown in Fig. 8, our model can generalize to handle text with diverse categories not seen at training such as umbrella, car, and computer.
Figure 7: Diversity: different 3D shapes generated with the same text prompt input.
Figure 8: Shapes generated with unseen categories.
We also provide some intriguing results generated by our methods. In Fig. 9, there are examples of some unusually shaped chairs and tables, that are not typically found in the real world but align well with their input text descriptions. This indicates that our method is not only effective in managing the correspondence task between text and shape during 3D generation but also possesses a notable level of creativity and the ability to generate diverse 3D content.
Figure 9: Visualization of creativity: some interesting but rarely seen 3D objects in the real world.
@@ 4.2.6 Ablation Studies
Although many parameters exist (e.g., Generator and NeRF), actually most of them are frozen as Fig. 2 demonstrates. As for the ICNNs [26], a typical structure is adopted. To evaluate the effectiveness of the major component $\mathcal{L}_{fide}=\mathbb{E}_{\mu_{S}}d(T_{f}(\boldsymbol{X}),\boldsymbol{Y})$ in the optimization objective function (16) of the Kantorovich potential $f$, we conduct the ablation study in two cases: (A) Without $\mathcal{L}_{fide}$; (B) With $\mathcal{L}_{fide}$. Tab. II summarizes the results of each case in terms of CLIP R-Precision and FID score. From the results, it can be seen that $\mathcal{L}_{fide}$ can significantly improve the performance of the generative model, especially in the CLIP R-Precision metric. The reason behind this is that $\mathcal{L}_{fide}$ provides one-to-one supervision of text and images. By comparing with our full objective function (bottom row), the necessity of $\mathcal{L}_{fide}$ can be validated.
Methods $L_{fide}$ CLIP R-Precision FID
A × 68.5 50.2
B 80.1 42.6
TABLE II: Ablation with regular term of objective function
@@ 5 Conclusion
In this paper, we propose a novel text-to-3D generation framework in combination with hyper-spherical optimal transport, i.e. HOTS3D, which can generate diverse and plausible 3D shapes from text prompts. Our method constructs the SOT map from the text feature hyper-sphere to the image feature hyper-sphere, achieving alignment between text and image features, while also compensating for the missing detailed information in high-level texts to some extent. The experimental results on synthetic data demonstrate the effectiveness of the proposed SOT algorithm. To address the text-to-3D generation, we utilized a pre-trained diffusion-based conditional generator and a controlled NeRF model to transform the aligned feature into 3D content. Our method is shown to generate 3D shapes with superior text semantic consistency compared to previous works, as demonstrated by extensive comparative experiments.
One limitation of our approach is the necessity for the pairing of text and image features to solve the SOT map, which limits the application scenarios of the algorithm. Additionally, optimizing two convex functions simultaneously during training can result in computational waste. In the future, we will study simplifying computational architecture.

### v2

@@ 4.2.3 Qualitative Results
The qualitative comparison on Text2Shape is presented in Fig. 6, indicating that baselines [16, 18, 7, 33] encounter challenges in generating precise and realistic 3D objects, resulting in semantic inconsistency or incomplete 3D shapes. The results of Kiss3DGen still have significant room for improvement in geometry and shape. Ln3Diff can produce 3D content with richer textures and colors, but struggles with semantic consistency. TAPS3D often produces simple tables and chairs, and Shap·E lacks geometric control, both needing better semantic alignment. Although SDFusion performs more reasonably, it still shows text-shape mismatches. Our method, integrating the SOT map, produces more coherent and higher-quality 3D objects, matching the text closely, as seen in examples like “2 tables hooked together”.
The qualitative comparison in Fig. 7 highlights results on Objaverse. Ln3diff produces outputs that deviate significantly from the input text. Kiss3DGen and Shap·E generate decent geometry and texture but struggle with semantic consistency. ProlificDreamer and RichDreamer, as optimization-based methods, achieve better semantic alignment but require several hours per prompt. In contrast, our method generates 3D models that align closely with the input text in an end-to-end fashion, while preserving higher quality and detail, demonstrating that semantic alignment can enhance the geometric quality of generated models.
Figure 7: Qualitative comparison on Objaverse. Red boxes on the results match the distinctive descriptions highlighted in red.
4.2.4 User studies
To enable a thorough comparison, we conducted a comprehensive user study focused on three key aspects: geometry quality, texture quality, and alignment with the text prompt. The study used a test set of 12 text prompts and involved 80 participants. As illustrated in Fig. 8, HOTS3D outperforms other methods in appearance, geometry, and semantic consistency for text-to-3D generation. Specifically, HOTS3D received 82.9% of the votes for geometry, 91.6% for texture, and 81.3% for semantic consistency. Considering the cost of participant involvement, we conducted the comparative experiments on a limited set of results (12 text prompts). Nonetheless, the observed user preferences still provide clear evidence of HOTS3D’s effectiveness.
Figure 8: User studies comparing HOTS3D with state-of-the-art methods show a strong preference for HOTS3D in terms of geometry quality, texture quality, and semantic consistency.
@@ 4.2.5 Ablation Studies
[…]
ViT-B/32 ViT-L/14
Xcube 0.790 0.796 3.0 0.150 0.953
Xcube+SOT 0.873 0.909 2.6 0.150 0.960
Ln3diff 0.835 0.862 2.8 0.139 0.954
Ln3diff+SOT 0.863 0.907 4.4 0.112 0.968
[…]
@@ 5 Conclusion
In this paper, we propose a novel text-to-3D generation framework in combination with hyper-spherical OT, i.e. HOTS3D, which can generate diverse and plausible 3D shapes from text prompts. Our method constructs the SOT map from the text feature hyper-sphere to the image feature hyper-sphere, achieving alignment between text and image features, while compensating to some extent for the lack of detailed information in high-concise text prompts. Then, a pre-trained diffusion-based conditional generator was utilized to transform the aligned feature into 3D content. Experimental results on synthetic data confirm the effectiveness of the proposed SOT algorithm for distribution alignment. Quantitative and qualitative comparisons on Text2Shape and Objaverse datasets demonstrate that our method can generate high-quality 3D shapes that accurately reflect text semantics. End-to-end distribution alignment allows the SOT module to be seamlessly integrated into other models, improving semantic consistency, while also making HOTS3D highly efficient with strong generalization and diversity.
One limitation of our approach is that it frequently generates inconsistent outputs when handling long texts. Another issue is that optimizing two convex functions simultaneously during training can lead to unnecessary computational overhead. In future work, we plan to simplify the computational architecture and explore more powerful vision language models to improve output consistency.